# Auriga Streams I: disrupting satellites surrounding Milky Way-mass haloes at multiple resolutions

### v1

@@ 6.4 The Auriga haloes in context
The first of these is that the stellar mass-halo mass relation in Auriga is considerably higher than in comparable cosmological simulations (see Figure 2 of Sales et al., 2022) and observational constraints (Behroozi et al., 2013; Moster et al., 2013), particularly in the stellar mass range of $10^{8}-10^{10}$ M (though see Posti et al. 2019; Grand et al. 2024 for discussion on the uncertainty in this relation and Simpson et al. 2018 for comparisons between the Auriga and Milky Way satellite luminosity functions). As an example, an Auriga central galaxy with $M_{\ast}\sim 10^{9}$ M can expect to live in a halo of $M_{\text{200c}}\sim 2\times 10^{10}$ M, while a galaxy of the same stellar mass in APOSTLE enjoys a much more spacious halo of $M_{\text{200c}}\sim 10^{11}$ M. This probably means that the number of objects at this stellar mass scale is slightly higher and such objects experience less dynamical friction in Auriga than in other simulations (since they would occupy more massive haloes, which are rarer in $\Lambda$CDM). Both of these effects boost the number of massive ($M_{\ast}>10^{8}$ M) streams.
@@
As an example, in Figure 12 we compare the stellar mass functions of streams in our sample at different resolution levels, as well as to those identified in FIRE (Panithanpaisal et al., 2021) which are similar in mass resolution to our level 3. The apparent disagreement between level 4 and 3 arises from the assembly histories of the level 3 haloes; the median of the level 4 runs for only Au-(6, 16, 21, 23, 24, 27), shown as a dashed line, agrees with the median for level 3. The apparent disagreement between level 2 and the other Auriga levels is due to the quiet merger history of Au-6; the lower resolution runs of the same halo have similarly low numbers of streams relative to the rest of the Auriga sample. Finally, at fixed stellar mass we find that there are approximately twice as many streams in Auriga as in FIRE, an effect which is not caused by differences in our definitions of a stream (see Appendix A). It is possible that a combination of different assembly histories and stellar mass-halo mass relations could alleviate the difference in stream mass functions between FIRE and Auriga, or it could still point towards interesting differences in galaxy formation physics (see Paper II for further discussion on this point). Stream mass functions in comparable suites of cosmological simulations of Milky Way-mass haloes might be illuminating on this point.
Finally, it is possible that numerical issues related to the finite mass resolution of cosmological simulations could affect how satellites disrupt. Recent studies using tailored $N$-body simulations have highlighted that subhaloes can artificially disrupt due to an insufficient number of particles or force resolution (van den Bosch & Ogiya, 2018; Errani & Peñarrubia, 2020; Errani & Navarro, 2021), though such experiments do not capture the important effects of preprocessing (He et al., 2024) or the tidal field of a central massive disk (Garrison-Kimmel et al., 2017; Kelley et al., 2019; Wang et al., 2024). In addition, spurious heating of galaxies due to unequal dark matter and star particle masses or insufficient total number of particles (Ludlow et al., 2019, 2020, 2021, 2023) could make satellites easier to disrupt into streams; the impact of spurious heating can be tested by repeating this analysis with forthcoming simulations that explicitly account for this effect (e.g. colibre; Schaye et al. in preparation).
@@ 6.5 Convergence across resolution
Figure 13: Convergence of accreted objects from Au-6 in the stellar mass-velocity dispersion plane used to classify objects as stream or phase-mixed (as in Figure 3). Each point corresponds to a system at a given resolution level, with dashed lines connecting objects to their counterpart one resolution level higher. Some objects only begin to be resolved at level 3, while objects that are only resolved in level 2 are not shown. Most accretion events are sufficiently converged within this plane (not crossing the classification boundary) or edge cases (fall near the boundary), lending confidence to our classification.
@@
We find that there is a slight systematic trend that satellites are better at retaining stars with increasing resolution. For the 65 systems matched from level 4 to 3 with an absolute change in $f_{\text{bound}}$ of at least 1 per cent [^12] We note that this particular statistic emphasises changes in $f_{\text{bound}}$ with resolution. Many objects consistently have $f_{\text{bound}}\sim 0$ or $\sim$1., this difference is $f_{\text{bound}}^{\text{high}}-f_{\text{bound}}^{\text{low}}=0.04^{+0.15}_{-0.08}$. From level 3 to 2, the 11 matched systems satisfying the same condition have $f_{\text{bound}}^{\text{high}}-f_{\text{bound}}^{\text{low}}=0.07^{+0.10}_{-0.17}$. In addition, 13 (7) accretions from level 4 to 3 (3 to 2) have no bound progenitor at lower resolution but some progenitor ($f_{\text{bound}}>0$) at higher resolution. While substantial changes (over $\sim$30 per cent) in $f_{\text{bound}}$ are often instances of different accretion timing and number of experienced pericentric passages (see Section 4), these results suggest some impact of numerical resolution on individual systems (van den Bosch & Ogiya, 2018; Errani & Peñarrubia, 2020). However, we note that the bulk of systems appear remarkably similar (Figure 6) and that global properties of all accreted stellar mass are largely preserved (Table 1 and Figure 9).
@@
In Figure 13 we show the distribution of accretions in the same $M_{\ast}-\sigma_{\text{vel}}^{50}$ plane that we use to classify systems as streams or phase-mixed, but now connecting objects matched in Au-6 at all resolution levels (similar trends apply to other haloes simulated at level 3). We only include objects that have a lower resolution counterpart, such that some systems are only represented in levels 3 and 2. Structures that are only resolved in level 2 are omitted. While most accretions (especially at higher mass) do not change substantially in this view, two systematic trends with resolution emerge. The first is that individual systems, particularly those with $M_{\ast}>10^{7}$ M, have higher stellar masses with increasing resolution, an effect that has been noted in previous Auriga studies (Grand et al., 2017, 2021). The second is that phase-mixed systems at the lower mass end of a given resolution level ($\lesssim 10^{7}$ M for level 4, $\lesssim 10^{6}$ M for level 3) tend to have substantially lower $\sigma_{\text{vel}}^{50}$ in their higher resolution counterparts. This is an apparent difference (not attributable to the Auriga model) due to our requirement that at least $k=7$ neighbours are used when calculating the local velocity dispersion at lower resolution; the change comes from switching to 1 per cent of star particles in the object at higher resolutions. We note that even with both of these effects, the variations for individual systems in this plane do not typically result in crossing the stream/phase-mixed separation line, lending confidence to our classification method.
@@ Acknowledgements
We thank Emily Cunningham, Denis Erkal, Danny Horta, Nondh Panithanpaisal, Robyn Sanderson, Eugene Vasiliev, and Anna Wright for enlightening discussions. AHR thanks Taylor Swift’s The Tortured Poets Department and Lindsey Stirling’s Duality for providing a soundscape condusive to writing this manuscript. This project was developed in part at the Streams24 meeting hosted at Durham University.
@@ Appendix B Catalogue of accreted structures
Halo Level ID Morphology $\log_{10}(M_{\ast}/\text{M}_{\odot})$ $N_{\text{part}}$ $f_{\text{bound}}$ Preprocessed Match ID Match quality

### v2

@@ 6.4 The Auriga haloes in context
[…]
Figure 13: Convergence of accreted objects in $f_{\text{bound}}$ and the stellar mass-velocity dispersion plane used for morphological classification. Left: change in $f_{\text{bound}}$ between objects matched across resolution levels, excluding objects that have no progenitor at both resolutions $(f_{\text{bound}}^{\text{high}}=f_{\text{bound}}^{\text{low}}=0)$. Dashed lines indicate an absolute change of $\pm 5$ per cent and the y-axis units are arbitrary to normalise both histograms to unity. Right: stellar mass-velocity dispersion plane, as presented in Figure 3, only showing systems in Au-6 for visual clarity. Each point corresponds to a system at a given resolution level, with dashed lines connecting objects to their counterpart one resolution level higher. Some objects only begin to be resolved at level 3, while objects that are only resolved in level 2 are not shown. Most accretion events are sufficiently converged within the stellar mass-velocity dispersion plane, lending confidence to our stream/phase-mixed classification, while we find a modest preferential increase in $f_{\text{bound}}$ with increasing resolution.
@@
As an example, in Figure 12 we compare the stellar mass functions of streams in our sample at different resolution levels, as well as to those identified in FIRE (Panithanpaisal et al., 2021) which are similar in mass resolution to our level 3. The apparent disagreement between level 4 and 3 arises from the assembly histories of the level 3 haloes; the median of the level 4 runs for only Au-(6, 16, 21, 23, 24, 27), shown as a dashed line, agrees with the median for level 3. The apparent disagreement between level 2 and the other Auriga levels is due to the quiet merger history of Au-6; the lower resolution runs of the same halo have similarly low numbers of streams relative to the rest of the Auriga sample (see Appendix C). Finally, at fixed stellar mass we find that there are approximately twice as many streams in Auriga as in FIRE, an effect which is not caused by differences in our definitions of a stream (see Appendix A). It is possible that a combination of different assembly histories and stellar mass-halo mass relations could alleviate the difference in stream mass functions between FIRE and Auriga, or it could still point towards interesting differences in galaxy formation physics (see Paper II for further discussion on this point). Stream mass functions in comparable suites of cosmological simulations of Milky Way-mass haloes might be illuminating on this point.
Finally, it is possible that numerical issues related to the finite mass resolution of cosmological simulations could affect how satellites disrupt. Recent studies using tailored $N$-body simulations have highlighted that subhaloes can artificially disrupt due to an insufficient number of particles or force resolution (van den Bosch & Ogiya, 2018; Errani & Peñarrubia, 2020; Errani & Navarro, 2021), though such experiments do not capture the important effects of preprocessing (He et al., 2025) or the tidal field of a central massive disk (Garrison-Kimmel et al., 2017; Kelley et al., 2019; Wang et al., 2025). In addition, spurious heating of galaxies due to unequal dark matter and star particle masses or insufficient total number of particles (Ludlow et al., 2019, 2020, 2021, 2023) could make satellites easier to disrupt into streams; the impact of spurious heating can be tested by repeating this analysis with forthcoming simulations that explicitly account for this effect (e.g. colibre; Schaye et al. in preparation).
@@ 6.5 Convergence across resolution
We find that there is a slight systematic trend that satellites are better at retaining stars with increasing resolution. In Figure 13 we show the change in $f_{\text{bound}}$ for objects that have a progenitor in either resolution level. For the 65 systems matched from level 4 to 3 with an absolute change in $f_{\text{bound}}$ of at least 1 per cent [^13] We note that this particular statistic emphasises changes in $f_{\text{bound}}$ with resolution. Many objects consistently have $f_{\text{bound}}=0$ or =1=1 (55 per cent from level 4 to 3 and 61 per cent from level 3 to 2)., this difference is $f_{\text{bound}}^{\text{high}}-f_{\text{bound}}^{\text{low}}=0.04^{+0.15}_{-0.08}$. From level 3 to 2, the 11 matched systems satisfying the same condition have $f_{\text{bound}}^{\text{high}}-f_{\text{bound}}^{\text{low}}=0.07^{+0.10}_{-0.17}$. In addition, 13 (7) accretions from level 4 to 3 (3 to 2) have no bound progenitor at lower resolution but some progenitor ($f_{\text{bound}}>0$) at higher resolution. Finally, we can assess the impact of resolution on actively disrupting satellites by considering objects that are classified as streams in both level 4 and level 3. Of these 60 streams, 50 (83 per cent) have $f_{\text{bound}}>0$ in level 4 and 59 (98 per cent) in level 3. The nine streams that gain a progenitor in level 3 have $f_{\text{bound}}$ between 0.0006 and 0.08, suggesting that they are in a similar dynamical state in level 4 but the progenitor is not recovered at lower resolution. While substantial changes (over $\sim$30 per cent) in $f_{\text{bound}}$ are often instances of different accretion timing and number of experienced pericentric passages (see Section 4), these results suggest some impact of numerical resolution on individual systems (van den Bosch & Ogiya, 2018; Errani & Peñarrubia, 2020). However, we note that the bulk of systems appear remarkably similar (Figure 6) and that global properties of all accreted stellar mass are largely preserved (Table 1 and Figure 9).
@@
In Figure 13 we also show the distribution of accretions in the same $M_{\ast}-\sigma_{\text{vel}}^{50}$ plane that we use to classify systems as streams or phase-mixed, but now connecting objects matched in Au-6 at all resolution levels (similar trends apply to other haloes simulated at level 3). We only include objects that have a lower resolution counterpart, such that some systems are only represented in levels 3 and 2. Structures that are only resolved in level 2 are omitted. While most accretions (especially at higher mass) do not change substantially in this view, two systematic trends with resolution emerge. The first is that individual systems, particularly those with $M_{\ast}>10^{7}$ M, have higher stellar masses with increasing resolution, an effect that has been noted in previous Auriga studies (Grand et al., 2017, 2021). The second is that phase-mixed systems at the lower mass end of a given resolution level ($\lesssim 10^{7}$ M for level 4, $\lesssim 10^{6}$ M for level 3) tend to have substantially lower $\sigma_{\text{vel}}^{50}$ in their higher resolution counterparts. This is an apparent difference (not attributable to the Auriga model) due to our requirement that at least $k=7$ neighbours are used when calculating the local velocity dispersion at lower resolution; the change comes from switching to 1 per cent of star particles in the object at higher resolutions. We note that even with both of these effects, the variations for individual systems in this plane do not typically result in crossing the stream/phase-mixed separation line, lending confidence to our classification method.
@@ Acknowledgements
We thank Emily Cunningham, Denis Erkal, Danny Horta, Nondh Panithanpaisal, Robyn Sanderson, Eugene Vasiliev, and Anna Wright for enlightening discussions, as well as an anonymous referee for helpful suggestions regarding its presentation. AHR thanks Taylor Swift’s The Tortured Poets Department and Lindsey Stirling’s Duality for providing a soundscape condusive to writing this manuscript. This project was developed in part at the Streams24 meeting hosted at Durham University.
@@ Appendix B Catalogue of accreted structures
Halo Level ID Morphology $\log_{10}(M_{\ast}/M_{\odot})$ $N_{\text{part}}$ $f_{\text{bound}}$ Preprocessed Match ID Match quality
@@
Appendix C Stream stellar mass functions for Au-6
Figure 16: Stellar mass functions for stellar streams, focusing only on Au-6 at different resolution levels (see Figure 12 for all Auriga haloes). The apparent disagreement between level 4 and the other levels at $M_{\ast}>10^{7}$ M and between levels 3 and 2 at $M_{\ast}<3\times 10^{6}$ M is due to matched objects having different classifications at different levels (see text of Appendix C for details).
In Figure 16 we present the stellar mass function for streams for Au-6 at various resolution levels, as a complement to Figure 12. At first glance, the apparent disagreement between level 4 and the other levels at $M_{\ast}>10^{7}$ M and between levels 3 and 2 at $M_{\ast}<3\times 10^{6}$ M suggest that these stream mass functions are not converged across resolution. However, this is typically due to individual accretion events having intact or phase-mixed classifications at other levels. In particular:
  • The most massive stream in level 4 is classified as intact in levels 3 and 2. This is the same system shown in the leftmost panel of Figure 6.
  • Three streams in level 3 are classified as phase-mixed in level 2. All three systems can be seen crossing the stream/phase-mixed boundary in Figure 13, while the most massive is the case highlighted in Figure 6 (third panel from the right) and Figure 7 (bottom panels). All level 2 streams with $M_{\ast}>4\times 10^{5}$ M are also streams in level 3, while objects below this mass do not appear in the level 3 catalogues.
These cases arise from looking at a single halo with a strict definition for ‘stream’, which will inevitably lead to boundary cases. Since Au-6 has a very quiet merger history with few mergers overall, variations are relatively impactful on the stream stellar mass function. These effects are mitigated when considering several haloes at once, as is the case in Figure 12.